EB011

PI: Barthel
Title: M 104

Session: Oct 1996 (session 3)
Frequency: 6 cm
Mode: B
Obs time (UT): 298 EU 07:00(24/10)-14:00(24/10)
Stations: EbJbNtOnWb
Processor: Bonn
EU Mk3 Tapes/St.: 1

Comment


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (tony foley at Thu Oct 24 06:01:21 EDT 1996) Minor failures
System clock died at 09:21UT, and repaired about 9:45UT. This meant that the telescope pointing was wrong from 9:21 to ~9:50UT

Onsala (A. Polatidis, J. Conway at Thu Oct 24 17:56:43 EDT 1996) Success
No known problems. Weather good, moderate winds.

Effelsberg (D.Graham at Mon Oct 28 10:39:06 EST 1996) Success
No known problems

Noto (Corrado Trigilio at Wed Oct 30 05:40:30 EST 1996) Success
No known problems. Clear sky.

Jodrell (Alastair Gunn at Thu Nov 7 06:43:18 EST 1996) Success
No known problems. Tape MEDI2001 shipped to Bonn on 28/10/96 via Wodocroft.

MPIfR (Scott Aaron at Tue Sep 23 14:23:11 METDST 1997) Minor failures
Westerbork 100% data loss--instrumental polarization > 150% Medicina did not participate